# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v2

@@ II.4 Perturbation Theory
The valley splitting can be approximated by first-order degenerate perturbation theory [15] assuming that both the intervalley-coupling potential $V_{c}\left(\mathbf{r}\right)$ and the alloy-disorder potential $\delta U_{\mathrm{\mathrm{het}}}\left(\mathbf{r}\right)$ are small. In this case, the problem (1) is to zeroth order approximated by
$\left(-\frac{\hbar^{2}}{2}\nabla\cdot\left(m^{-1}\nabla\right)+\left\langle U\left(\mathbf{r}\right)\right\rangle\right)\Psi_{0}\left(\mathbf{r}\right)$ $=E_{0}\Psi_{0}\left(\mathbf{r}\right),$
where $\left\langle U\left(\mathbf{r}\right)\right\rangle=U_{\mathrm{QD}}\left(x,y\right)+U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)$ is the mean potential energy. The reduced zeroth-order problem describes two energetically degenerate valley states with identical orbital wave function $\Psi_{0}\left(\mathbf{r}\right)$. Using the separation ansatz $\Psi_{0}\left(\mathbf{r}\right)=\phi_{0}\left(x,y\right)\psi_{0}\left(z\right)$ and $E_{0}=E_{t,0}+E_{l,0}$, the problem decouples into two scalar effective mass-type Schrödinger equations, i.e., the transverse problem
@@
with ground state energy $E_{t,0}=\left(\hbar\text{$\omega$}_{x}+\hbar\text{$\omega$}_{y}\right)/2$ and QD width $l_{j}=\sqrt{\hbar/\left(m_{t}\omega_{j}\right)}$, $j\in\left\{x,y\right\}$. The ground state $\left\{E_{l,0},\psi_{0}\left(z\right)\right\}$ of the longitudinal problem is obtained numerically using a finite difference approximation.
In first-order perturbation theory, the multi-valley envelope function in Eq. (1) for the degenerate ground state subspace is taken as $\left(\Psi_{+}\left(\mathbf{r}\right),\Psi_{-}\left(\mathbf{r}\right)\right)^{T}\approx\left(\eta_{+},\eta_{-}\right)^{T}\Psi_{0}\left(\mathbf{r}\right)$, which after integration with $\Psi_{0}^{*}\left(\mathbf{r}\right)$ yields the reduced problem
$\left(\begin{array}[]{cc}E_{0}+\delta E_{0}&\Delta\\ \Delta^{*}&E_{0}+\delta E_{0}\end{array}\right)\left(\begin{array}[]{c}\eta_{+}\\ \eta_{-}\end{array}\right)=E\left(\begin{array}[]{c}\eta_{+}\\ \eta_{-}\end{array}\right).$ (19)
Here, $\delta E_{0}=\int\mathrm{d}^{3}r\,\delta U_{\mathrm{het}}\left(\mathbf{r}\right)\Psi_{0}^{2}\left(\mathbf{r}\right)$ is a random shift of the energy level (that has no further physical implications) and the parameter
$\Delta$ $=\int\mathrm{d}^{3}r\,\Psi_{0}^{*}\left(\mathbf{r}\right)V_{c}\left(\mathbf{r}\right)\Psi_{0}\left(\mathbf{r}\right)$
$=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\times$ (20)
×$\hphantom{=}\times\int\mathrm{d}^{3}r\,\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{r}}U\left(\mathbf{r}\right)\Psi^{2}_{0}\left(\mathbf{r}\right)$
is the complex-valued intervalley-coupling strength. We note that Eq. (20) involves all of the aforementioned effects, i.e., strain (via modification of the reciprocal lattice vectors and the Bloch factor expansion coefficients), alloy disorder and non-trivial resonances due to coupling of valley states within different Brillouin zones. Finally, the energy splitting of the two valley states described by the reduced problem (19) is obtained as
@@
A comprehensive theoretical description of the intervalley-coupling parameter provides the key to the engineering of deterministic enhancements of the valley splitting in Si/SiGe qubits.
@@ II.5 Intervalley Coupling Parameter
The intervalley coupling parameter $\Delta$ given in Eq. (20) has a deterministic and a random component
@@
reflecting the deterministic and stochastic components of the total confinement potential (4). In the following, we will evaluate both components.
@@ II.5.1 Deterministic Component
The deterministic component
$\Delta_{\mathrm{det}}=\Delta_{\mathrm{det}}^{l}+\Delta_{\mathrm{det}}^{t}$ (23)
comprises a transverse and a longitudinal contribution corresponding to the confinement potentials of the two decoupled effective mass equations (16)–(17). Using the exact ground state wave function (18) of the transverse problem, the in-plane integration can be evaluated explicitly. With this, one arrives at
$\Delta_{\mathrm{det}}^{l}$ $=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\times$ (24)
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(G_{z}-G_{z}^{\prime}+2k_{0}\right)z}\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi_{0}^{2}\left(z\right)$
and
$\Delta_{\mathrm{det}}^{t}$ $=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\times$ (25)
×$\phantom{=}\times\bigg{[}\frac{\hbar\omega_{x}}{2}\left(\frac{1}{2}-\frac{\left(G_{x}-G_{x}^{\prime}\right)^{2}l_{x}^{2}}{4}\right)$
$+\frac{\hbar\omega_{y}}{2}\left(\frac{1}{2}-\frac{\left(G_{y}-G_{y}^{\prime}\right)^{2}l_{y}^{2}}{4}\right)\bigg{]}\times$
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(G_{z}-G_{z}^{\prime}+2k_{0}\right)z}\psi_{0}^{2}\left(z\right).$
@@
$S\left(z\right)=\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi^{2}_{0}\left(z\right)$
@@
gives a strong contribution via an enhancement of the coupling strength between certain valley states, possibly across different Brillouin zones. A similar reasoning holds true for an enhancement of the transverse component (25) with correspondingly adapted $S\left(z\right)$, but for typical parameter values the longitudinal component (24) is dominant. This concept of confinement potential engineering to achieve large Fourier amplitudes $\tilde{S}\left(q\right)$ is explicitly addressed in the wiggle well proposal [34, 31, 33], but is implicitly exploited also in other approaches (e.g., sharp interfaces [24], Ge-spike [35]).
The two expressions (24) and (25) can be considerably simplified by exploiting the fact that the in-plane extension of the QD wave function is much larger than the lattice constant $l_{x},l_{y}\gg a_{0}$. With this, the Gaussians (second line of both expressions), effectively reduce to a Kronecker-Delta reproducing the selection rule for quantum wells [16]. In the limit of small strain, this approximation, which is described in detail in Appendix B, allows for a very compact notation
@@
where we have introduced the effective band structure coefficients $C_{n}^{\left(2\right)}$ described in Eq. (58) and the integrals $J_{n}^{\mathrm{det}}$ defined in Eq. (68). The compact form allows for a fast numerical evaluation, where a single summation over a few integers replaces the tedious double summation over the reciprocal lattice vectors in Eqs. (24)–(25).
Fig. 2: (a) Band structure coefficients $C_{n}^{\left(2\right)}$ governing the magnitude of the deterministic contribution to the valley splitting as a function of shear strain $\varepsilon_{x,y}$. The coefficients have been computed from Eq. (58) using plane wave expansions of the Bloch factors for strained silicon at the conduction band minimum. In addition to shear strain, biaxial tensile strain arising from the $\mathrm{Si}/\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ heterostructure was assumed, see Eq. (15). Coefficients with odd $n$ show a linear dependence on shear strain, while coefficients with even $n$ are practically constant. (b) Same as (a) for the band structure coefficients $C_{n}^{\left(4\right)}$ given by Eq. (67) governing the magnitude of the disorder-induced contribution to the valley splitting.
@@
is taken into account. The plot of the band structure coefficients in Fig. 2 (a) shows, that this approximation is indeed justified in many scenarios, but might fail when the integrands of Eqs. (24)–(25) exhibit strong non-trivial resonances, i.e., with $\mathbf{G}-\mathbf{G}^{\prime}=n\mathbf{G}_{0}$ for $n\neq 0$ and $\mathbf{G}_{0}$ given in Eq. (62). In these cases, contributions proportional to band structure coefficients $C^{\left(2\right)}_{n\neq 0}$ become relevant. A particularly important resonance is the one mediated by $C_{n=-1}^{\left(2\right)}$ at
@@
which is the reciprocal-space distance of the valley states from the Brillouin zone boundary, see Fig. 1 (a). We note that uniaxial strain $\varepsilon_{z,z}$ (which is typically compressive) leads to a slight modification of this distance. As shown in Fig. 2, the band structure coefficient $C_{n=-1}^{\left(2\right)}$ obtained from the EPM is linearly dependent on the shear strain component as $C_{n=-1}^{\left(2\right)}\approx-0.589\times\varepsilon_{x,y}$. This result is in good agreement with the theory recently presented by Woods et al. [33], which predicts a linear dependency of the long-period wiggle-well on shear strain following a $\mathrm{s}\mathrm{p}^{3}\mathrm{d}^{5}\mathrm{s}^{*}$ tight-binding model. One can clearly see that the resonance is suppressed in the absence of shear strain $\varepsilon_{x,y}=0$, which is explained by a nonsymmorphic screw symmetry of the (either relaxed or biaxially strained) silicon crystal structure [31, 33]. In the presence of shear strain, however, this symmetry is broken such that the resonance at $n=-1$ yields enhancements also for other strategies much simpler than the long-period wiggle-well, e.g., sharp interfaces or uniform Ge concentrations in the QW. This is discussed in more detail along with numerical results in Sec. III below. Finally, in Sec. III we also observe small contributions from the non-trivial $n=-2$ resonance, in particular at sharp interfaces. This corresponds to
@@ II.5.2 Random Component
×$\phantom{=}\times\int\mathrm{d}^{3}r\,\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{r}}\,\delta U_{\mathrm{het}}\left(\mathbf{r}\right)\Psi_{0}^{2}\left(\mathbf{r}\right).$
As shown in Appendix C.1, $\Delta_{\mathrm{rand}}$ obeys a complex normal distribution
@@
with zero mean $\mu=0$, covariance $\Gamma=\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$ and relation (or pseudo-covariance) parameter $C=\langle\Delta_{\mathrm{rand}}^{2}\rangle$. As the relation parameter is typically negligible in comparison to the covariance, the random contribution is well approximated by a circular symmetric normal distribution in the complex plane with independent and identically distributed real and imaginary parts
$\mathrm{Re}\left(\Delta_{\mathrm{rand}}\right)$ $\sim\mathrm{Normal}\left(\mu=0,\sigma^{2}=\frac{1}{2}\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle\right),$
$\mathrm{Im}\left(\Delta_{\mathrm{rand}}\right)$ $\sim\mathrm{Normal}\left(\mu=0,\sigma^{2}=\frac{1}{2}\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle\right).$
We refer to Appendix C.1 for a detailed derivation. Consequently, the characterization of the disorder-induced component to the valley splitting requires solely the computation of the covariance $\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$. Using the covariance function of the random potential (14) and the in-plane wave functions (18), one obtains
$\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle=\frac{1}{2\pi l_{x}l_{y}}\left(\Delta E_{c}\right)^{2}\Omega_{a}\times$ (34)
×$\phantom{==}\times\sum_{\mathbf{G},\mathbf{G}^{\prime}}\sum_{\mathbf{K},\mathbf{K}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)c_{+}\left(\mathbf{K}\right)c_{-}^{*}\left(\mathbf{K}^{\prime}\right)\times$
×$\phantom{==}\times\mathrm{e}^{-\frac{1}{2}\big{(}\frac{(G_{x}-G_{x}^{\prime}-K_{x}+K_{x}^{\prime})l_{x}}{2}\big{)}^{2}}\mathrm{e}^{-\frac{1}{2}\big{(}\frac{(G_{y}-G_{y}^{\prime}-K_{y}+K_{y}^{\prime})l_{y}}{2}\big{)}^{2}}\times$
×$\phantom{==}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(G_{z}-G_{z}^{\prime}-K_{z}+K_{z}^{\prime}\right)z}X\left(z\right)\left(1-X\left(z\right)\right)\psi_{0}^{4}\left(z\right).$
@@
where the integrals $J_{n}^{\mathrm{rand}}$ are given in Eq. (66) and the band structure coefficients $C_{n}^{\left(4\right)}$ are defined in Eq. (67). We refer to Appendix B for details. The band structure coefficients $C_{n}^{\left(4\right)}$ are shown as a function of shear strain in Fig. 2 (b). In contrast to the deterministic part, non-trivial resonances are less pronounced in the disorder-induced contribution to the valley splitting such that Eq. (35) is in fact largely dominated by the trivial contribution proportional to $C_{n=0}^{\left(4\right)}$. In general, modifications due to shear strain are weaker for the random component than for the deterministic part.

### v3

@@ II.4 Perturbation Theory
The valley splitting can be approximated using first-order degenerate perturbation theory by assuming that both the intervalley-coupling term $V_{c}\left(\mathbf{r}\right)$ and the disorder potential $\delta U_{\mathrm{\mathrm{het}}}\left(\mathbf{r}\right)$ can be treated as small perturbations.
II.4.1 Unperturbed Problem
The unperturbed problem corresponding to Eq. (1) is the single-valley Schrödinger equation
$E_{0}\Psi_{0}\left(\mathbf{r}\right)$ $=-\frac{\hbar^{2}}{2}\nabla\cdot\left(m^{-1}\nabla\Psi_{0}\left(\mathbf{r}\right)\right)$ (16)
$\phantom{=}+\left(U_{\mathrm{QD}}\left(x,y\right)+U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\Psi_{0}\left(\mathbf{r}\right)$
where $\left\langle U\left(\mathbf{r}\right)\right\rangle=U_{\mathrm{QD}}\left(x,y\right)+U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)$ is the mean potential energy and $m=\mathrm{diag}\left(m_{t},m_{t},m_{l}\right)$ is the effective mass tensor. The unperturbed problem (16) describes two energetically degenerate (decoupled) valley states with identical orbital wave function $\Psi_{0}\left(\mathbf{r}\right)$. Using the separation ansatz $\Psi_{0}\left(\mathbf{r}\right)=\phi_{0}\left(x,y\right)\psi_{0}\left(z\right)$ and $E_{0}=E_{t,0}+E_{l,0}$, the problem separates into two scalar effective mass-type Schrödinger equations, i.e., the transverse problem
@@ II.4.1 Unperturbed Problem
with ground state energy $E_{t,0}=\left(\hbar\text{$\omega$}_{x}+\hbar\text{$\omega$}_{y}\right)/2$ and QD width $l_{j}=\sqrt{\hbar/\left(m_{t}\omega_{j}\right)}$, $j\in\left\{x,y\right\}$. The ground state $\left\{E_{l,0},\psi_{0}\left(z\right)\right\}$ of the longitudinal problem is computed numerically using a finite difference approximation. As the unperturbed single-valley problem (16) is identical for both valleys, the ground state is twofold degenerate.
II.4.2 First-Order Degenerate Perturbation Theory
A perturbative expression for the valley splitting is obtained using first-order degenerate perturbation theory [15]. The degenerate ground state subspace is spanned by the two independent basis states $\boldsymbol{\Psi}_{+}\left(\mathbf{r}\right)=\left(\Psi_{0}\left(\mathbf{r}\right),0\right)^{T}$ and $\boldsymbol{\Psi}_{-}\left(\mathbf{r}\right)=\left(0,\Psi_{0}\left(\mathbf{r}\right)\right)^{T}$. Substitution of the linear combination $\eta_{+}\boldsymbol{\Psi}_{+}\left(\mathbf{r}\right)+\eta_{-}\boldsymbol{\Psi}_{-}\left(\mathbf{r}\right)=\left(\eta_{+},\eta_{-}\right)^{T}\Psi_{0}\left(\mathbf{r}\right)$ into the equation for the first-order correction yields—after projection on the basis states—an eigenvalue equation for the first-order energy correction $E_{1}$
$\left(\begin{array}[]{cc}\delta E&\Delta\\ \Delta^{*}&\delta E\end{array}\right)\left(\begin{array}[]{c}\eta_{+}\\ \eta_{-}\end{array}\right)=E_{1}\left(\begin{array}[]{c}\eta_{+}\\ \eta_{-}\end{array}\right).$
Here we introduced the complex-valued intervalley-coupling parameter
$\Delta$ $=\int\mathrm{d}^{3}r\,\Psi_{0}^{*}\left(\mathbf{r}\right)V_{c}\left(\mathbf{r}\right)\Psi_{0}\left(\mathbf{r}\right)$ (20)
$=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\times$
[…]
and the disorder-induced energy shift
$\delta E=\int\mathrm{d}^{3}r\,\Psi_{0}^{*}\left(\mathbf{r}\right)\delta U_{\mathrm{het}}\left(\mathbf{r}\right)\Psi_{0}\left(\mathbf{r}\right).$ (21)
We note that Eq. (20) involves all of the aforementioned effects, i.e., strain (via modification of the reciprocal lattice vectors and the Bloch factor expansion coefficients), alloy disorder and non-trivial resonances due to coupling of valley states within different Brillouin zones. Finally, the energy correction $E_{1,\pm}=\delta E\pm\left|\Delta\right|$ yields the valley splitting
@@ II.4.2 First-Order Degenerate Perturbation Theory
An accurate description of the intervalley coupling parameter provides the key to the engineering of deterministic enhancements of the valley splitting in Si/SiGe qubits.
@@ II.5 Intervalley Coupling Parameter
The intervalley coupling parameter $\Delta$ has a deterministic and a random component
@@
reflecting the deterministic and the stochastic components of the total confinement potential (4):
$\Delta_{\mathrm{det}}$ $=\int_{V}\mathrm{d}^{3}r\,\mathrm{e}^{-2i\mathbf{k}_{0}\cdot\mathbf{r}}u_{+}^{*}\left(\mathbf{r}\right)u_{-}\left(\mathbf{r}\right)\left\langle U\left(\mathbf{r}\right)\right\rangle\left|\Psi_{0}\left(\mathbf{r}\right)\right|^{2},$
$\Delta_{\mathrm{rand}}$ $=\int_{V}\mathrm{d}^{3}r\,\mathrm{e}^{-2i\mathbf{k}_{0}\cdot\mathbf{r}}u_{+}^{*}\left(\mathbf{r}\right)u_{-}\left(\mathbf{r}\right)\delta U_{\mathrm{het}}\left(\mathbf{r}\right)\left|\Psi_{0}\left(\mathbf{r}\right)\right|^{2}.$
In the following, we will evaluate both components.
@@ II.5.1 Deterministic Component
Using the exact ground state wave function (19), the deterministic component of the intervalley coupling parameter is obtained after in-plane integration as
$\Delta_{\mathrm{det}}$ $=\sum_{\mathbf{G},\mathbf{G}^{\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)\times$ (24)
@@
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(G_{z}-G_{z}^{\prime}+2k_{0}\right)z}\,\bigg{[}U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)$
$\phantom{\times=}+\frac{\hbar\omega_{x}}{2}\left(\frac{1}{2}-\left(\frac{\left(G_{x}-G_{x}^{\prime}\right)l_{x}}{2}\right)^{2}\right)$
$\phantom{\times=}+\frac{\hbar\omega_{y}}{2}\left(\frac{1}{2}-\left(\frac{\left(G_{y}-G_{y}^{\prime}\right)l_{y}}{2}\right)^{2}\right)\bigg{]}\left|\psi_{0}\left(z\right)\right|^{2}.$
The Gaussian damping terms in the second line result from the shape of the in-plane wave function (19). The contributions from the in-plane QD potential (last two lines) are typically small.
@@
$S\left(z\right)=\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\left|\psi_{0}\left(z\right)\right|^{2}$
@@
gives a strong contribution via an enhancement of the coupling strength between certain valley states, possibly across different Brillouin zones. This concept of confinement potential engineering to achieve large Fourier amplitudes $\tilde{S}\left(q\right)$ is explicitly addressed in the wiggle well [34, 31, 33], but is implicitly employed also in other approaches (e.g., sharp interfaces [24], Ge-spike [35]).
The expression (24) can be considerably simplified by exploiting the fact that the in-plane extension of the QD wave function is much larger than the lattice constant $l_{x},l_{y}\gg a_{0}$. With this, the Gaussians in the second line of Eq. (24) effectively reduce to a Kronecker-Delta reproducing the selection rule for quantum wells [16], see Appendix C for details. This approximation allows for a very compact notation
@@
where we have introduced the coefficients $C_{n}^{\left(2\right)}$ described in Eq. (50) and the integrals $J_{n}^{\mathrm{det}}$ defined in Eq. (49).
Fig. 2: (a) Plot of the coefficients $C_{n}^{\left(2\right)}$ governing the magnitude of the deterministic contribution to the valley splitting as a function of shear strain $\varepsilon_{x,y}$. The coefficients have been computed from Eq. (50) using the plane wave expansion coefficients of the Bloch factors for strained silicon at the conduction band minimum. In addition to shear strain, biaxial tensile strain arising from the $\mathrm{Si}/\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ heterostructure was assumed, see Eq. (15). Coefficients with odd $n$ show a linear dependence on shear strain, while coefficients with even $n$ are practically constant. (b) Same as (a) for the coefficients $C_{n}^{\left(4\right)}$ given by Eq. (57) governing the magnitude of the disorder-induced contribution to the valley splitting.
@@
is taken into account. The plot of the coefficients $C_{n}^{\left(2\right)}$ in Fig. 2 (a) shows, that this approximation is indeed justified in many scenarios, but might fail when the integrand of Eq. (24) exhibits strong non-trivial resonances with valley states outside of the first Brillouin zone, i.e., with $\mathbf{G}-\mathbf{G}^{\prime}=n\mathbf{G}_{0}$ for $n\neq 0$ and $\mathbf{G}_{0}$ given in Eq. (54). In these cases, contributions proportional to other coefficients $C_{n\neq 0}^{\left(2\right)}$ become relevant. A particularly important resonance is the one at
@@
which is the reciprocal-space distance of the conduction band minimum from the Brillouin zone boundary, see Fig. 1 (a). We note that uniaxial strain $\varepsilon_{z,z}$ (which is typically compressive) leads to a slight modification of this distance. As shown in Fig. 2, the coefficient $C_{n=-1}^{\left(2\right)}$ obtained from the EPM is linearly dependent on the shear strain component as $C_{n=-1}^{\left(2\right)}\approx-32.0\times\varepsilon_{x,y}$. This result is in good agreement with the theory recently presented by Woods et al. [33], which predicts a linear dependency of the long-period wiggle-well on shear strain following a $\mathrm{s}\mathrm{p}^{3}\mathrm{d}^{5}\mathrm{s}^{*}$ tight-binding model. One can clearly see that the resonance is suppressed in the absence of shear strain $\varepsilon_{x,y}=0$, which is explained by a nonsymmorphic symmetry of the (either relaxed or biaxially strained) silicon crystal structure [31, 33]. In the presence of shear strain, however, this symmetry is broken such that the resonance at $n=-1$ also yields enhancements to structures such as sharp interfaces or QWs with uniform Ge concentrations, in addition to the long period wiggle well. This is discussed in more detail along with numerical results in Sec. III below. Finally, in Sec. III we will also observe small contributions from the $n=-2$ resonance, in particular at sharp interfaces. This corresponds to
@@ II.5.2 Random Component
×$\phantom{=}\times\int\mathrm{d}^{3}r\,\mathrm{e}^{-i\left(\mathbf{G}-\mathbf{G}^{\prime}+2\mathbf{k}_{0}\right)\cdot\mathbf{r}}\,\delta U_{\mathrm{het}}\left(\mathbf{r}\right)\left|\Psi_{0}\left(\mathbf{r}\right)\right|^{2}.$
As shown in Appendix D.1, $\Delta_{\mathrm{rand}}$ obeys a complex normal distribution
@@
with zero mean $\mu=0$, covariance $\Gamma=\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$ and pseudo-covariance $C=\langle\Delta_{\mathrm{rand}}^{2}\rangle$. As the pseudo-covariance is typically negligible in comparison to the covariance, the random contribution is well approximated by a circular symmetric normal distribution in the complex plane with independent and identically distributed real and imaginary parts
$\mathrm{Re}\left(\Delta_{\mathrm{rand}}\right)$ $\sim\mathrm{Normal}\left(\mu=0,\sigma^{2}=\frac{1}{2}\Gamma\right),$
$\mathrm{Im}\left(\Delta_{\mathrm{rand}}\right)$ $\sim\mathrm{Normal}\left(\mu=0,\sigma^{2}=\frac{1}{2}\Gamma\right).$
We refer to Appendix D.1 for a detailed derivation. Consequently, the characterization of the disorder-induced contribution to the valley splitting requires solely the computation of the covariance $\Gamma$. Using the covariance function of the random potential (14) and the in-plane wave functions (19), one obtains
$\Gamma=\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle=\frac{1}{2\pi l_{x}l_{y}}\left(\Delta E_{c}\right)^{2}\Omega_{a}$ (33)
×$\phantom{=}\times\sum_{\mathbf{G},\mathbf{G}^{\prime},\mathbf{G}^{\prime\prime},\mathbf{G}^{\prime\prime\prime}}c_{+}^{*}\left(\mathbf{G}\right)c_{-}\left(\mathbf{G}^{\prime}\right)c_{+}\left(\mathbf{G}^{\prime\prime}\right)c_{-}^{*}\left(\mathbf{G}^{\prime\prime\prime}\right)$
×$\phantom{=}\times\mathrm{e}^{-\frac{1}{2}\big{(}\frac{(G_{x}-G_{x}^{\prime}-G_{x}^{\prime\prime}+G_{x}^{\prime\prime\prime})l_{x}}{2}\big{)}^{2}}\mathrm{e}^{-\frac{1}{2}\big{(}\frac{(G_{y}-G_{y}^{\prime}-G_{y}^{\prime\prime}+G_{y}^{\prime\prime\prime})l_{y}}{2}\big{)}^{2}}$
×$\phantom{=}\times\int\mathrm{d}z\,\mathrm{e}^{-i\left(G_{z}-G_{z}^{\prime}-G_{z}^{\prime\prime}+G_{z}^{\prime\prime\prime}\right)z}X(z)\left(1-X(z)\right)\left|\psi_{0}(z)\right|^{4}.$
@@
where the integrals $J_{n}^{\mathrm{rand}}$ are given in Eq. (56) and the coefficients $C_{n}^{\left(4\right)}$ are defined in Eq. (57). We refer to Appendix C for details. The coefficients $C_{n}^{\left(4\right)}$ are plotted as a function of shear strain in Fig. 2 (b).